OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 8?

OpenStudy (anonymous):

the 2s are exponents

OpenStudy (anonymous):

The standard form of circle is (X-A)^2 + (Y-B)^2=R^2 Here if the given equation is compared with the equation of circle, then we get that R^2=8 Therefore R=Sq. root of 8 =\[\sqrt{8}=2\sqrt{2}\] is the radius :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're welcome. :)